# algorithmic

### v2

@@ 3.3 Enhancing Sample Generation with Noise Level Correction
The trained noise level correction network $r_{\theta}$ can be integrated into various existing sampling algorithms to improve sample quality. For algorithms that include an initial sample estimate $x_{0|t}$, eq. 5, we reformulate the one-step estimation as follows:
@@
We normalize the noise vector $\epsilon_{\theta}(\cdot)$ during the sampling as in eq. 17, process to decouple noise level (magnitude) correction $\hat{\sigma}_{t}$ from direction estimation $\hat{\epsilon}_{t}$. Empirical experiments show that normalizing $\epsilon_{\theta}(\cdot)$ and using $\hat{\sigma}_{t}$ to account for magnitude yields better results compared to not normalizing $\epsilon_{\theta}(\cdot)$. In the training loss function eq. 14, normalization of the noise vector $\epsilon$ is unnecessary because the randomly sampled noise $\epsilon$ naturally concentrates around the norm $\sqrt{n}$. However, the neural network-estimated noise vector $\epsilon_{\theta}(\cdot)$ does not maintain a constant norm.
@@
0:   Noise scheduler $\sigma_{t}$, randomness scale $\eta$ ($\eta=0$ for deterministic DDIM and $\eta=1$ for DDPM )
@@ 3.4 Constrained Sample Generation
Considering diffusion models, Theorem 1 demonstrates that the projection operator $\text{proj}_{\mathcal{K}}(\cdot)$ can be approximately computed using the denoising operator. Starting from an initial random point $x_{0}=\sigma_{\text{max}}\epsilon$, the projection onto the manifold $\mathcal{K}$ can be iteratively refined using eq. 16 and eq. 17, as follows:
@@
For the additional constraint projection, $\text{proj}_{\mathcal{C}}(x)$, the specific calculation depends on the nature of the constraint. The constraints for many image restoration tasks are linear, including inpainting, colorization, super-resolution, deblurring, and compressed sensing. For tasks with linear constraints, the projection can be computed directly or optimized using gradient descent. Consider an image restoration task formulated as $y=\textbf{A}x_{0}$, where $x_{0}$ represents the ground-truth image, $y$ is the degraded observation, A is the linear degradation operator. Given the degraded image $y$ and the current estimate $\hat{x}_{0|k}$, the projection onto the constraint can be computed as:
@@
Where $\textbf{A}^{\dagger}$ is the pseudo-inverse of A. In this work, we adopt the values of $\textbf{A}^{\dagger}$ for image restoration tasks as provided in Wang et al. (2023). Here, ${x}_{0|k}$ is the $k-$th th estimate satisfying $x_{0}\in\mathcal{K}\cap\mathcal{C}$. This iterative estimation requires a predefined noise scheduler $\sigma_{1},\cdots,\sigma_{(k)},\cdots$ to generate ${x}_{0|k}$ and $x_{(k+1)}$. To allow flexible and potentially unlimited refinement steps, we define the noise schedule $\sigma_{(k)}$ with a maximum noise level $\sigma_{(0)}=\sigma_{max}$ and a minimum level $\sigma_{min}$, decaying by a predefined factor $\eta<1$. If the noise level reaches $\sigma_{min}$, the process can either stop with returning ${x}_{0|k}$ or restart from $\sigma_{restart}$. This strategy permits an arbitrary number of refinement steps, stopping either at a desired loss threshold or continuing indefinitely. Since $\sigma_{(k)}$ represents the distance of noisy samples from the manifold, this decaying schedule incrementally reduces $x_{(k)}$’s distance from the manifold.
@@ 4.1 Toy Experiments
We conducted a toy experiment to demonstrate the effectiveness of noise level correction in sample generation for diffusion models. The objective was to generate samples on $d$-sphere manifold. The toy training dataset was sampled from $d$-dimensional sphere manifold embedded within an $n$-dimensional data space, where $d<n$. Detailed experimental design information is available in Section C.1.
After training, we applied the proposed 10-step DDIM with Noise Level Correction (DDIM-NLC), as detailed in Algorithm 1 to generate samples. This method was compared to the 10-step DDIM baseline. Our evaluation metric measured the distance between the generated samples and the ground-truth $d$-sphere manifold $\mathcal{K}$ (where lower distance indicates better results). Figure 4a presents the sample quality (measured as the distance to the manifold) for different methods across each denoising step. As shown, DDIM-NLC outperforms the baseline DDIM by generating samples that are consistently closer to the target manifold.
As we discussed in Theorom 1, the distance to the manifold can be approximated by the noise level, as shown in LABEL:eq:theorom_dist. We further evaluated the accuracy of the noise level correction by examining its proximity to the actual distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$. Figure 4b displays the relative distance estimation bias during sampling, calculated as
@@
Where $\hat{\sigma}_{t}=\sigma_{t}$ for DDIM, and $\hat{\sigma}_{t}=\sigma_{t}[1+r_{\theta}(\hat{x}_{t},\sigma_{t})]$ for DDIM-NLC. The results show that DDIM-NLC achieves a significantly smaller distance estimation bias than DDIM, particularly in the later steps as samples approach the manifold. The results of constrained sample generation are shown in Section C.2.
(a) Sample Quality
@@
Figure 4: Results from experiments on toy models. (a) Generated sample quality was evaluated the distance to the manifold. (b) Distance Estimation Error.
@@ 4.2 Unconstrained Image Generation
We conducted experiments to evaluate the effectiveness of noise level correction in unconstrained image generation tasks. The noise level correction network was trained on top of a pre-trained denoiser network. Notably, the noise level correction network is approximately ten times smaller than the denoiser network. Additional details on the experimental setup are provided in Section D.1. We used the FID (Fréchet Inception Distance) score as the evaluation metric to assess the quality of generated samples, where lower scores indicate better quality, following standard practice in image generation tasks Heusel et al. (2017). The experimental results for the DDIM/DDPM framework on the CIFAR-10 dataset are shown in Table 2. As observed, DDPM-NLC and DDIM-NLC, which incorporate noise level correction, outperform the original DDPM and DDIM models across all sampling steps. Specifically, our proposed noise level correction approach improves DDIM performance by 32%, 31%, 22%, 33%, and 28% for 10, 20, 50, 100, and 300 sampling steps, respectively.
@@ 4.3 Image Restoration
In this section, we evaluate the effectiveness of noise level correction on five common image restoration tasks, 4×4$\times$4 × super-resolution (SR) using bicubic downsampling, deblurring with a Gaussian blur kernel, colorization using an average grayscale operator, compressed sensing (CS) with a Walsh-Hadamard sampling matrix at a 0.25 compression ratio, and inpainting with text masks. These experiments are conducted on the ImageNet (Deng et al., 2009) and CelebA-HQ (Karras et al., 2018) datasets. We compare our method with recent state-of-the-art diffusion-based image restoration methods, including ILVR (Choi et al., 2021), RePaint plugmayr2022repaint, DDRM (Kawar et al., 2022), and DDNM (Wang et al., 2023). For a fair comparison, all diffusion-based methods utilize the same pretrained denoising networks with the same 100-step denoising process (100 number of inference steps), following the experimental setup in Wang et al. (2023). To evaluate sample quality, we use FID, PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index Measure). For colorization, where PSNR and SSIM are less effective metrics (Wang et al., 2023), we additionally use a Consistency metric, denoted as "Cons" and calculated as $\|Ax_{0}-y\|_{1}$. As a baseline, we also include the inverse solution for each image restoration task, given by $\hat{x}=\textbf{A}^{\dagger}y$, which achieves zero constraint violation but lacks the data manifold information.
@@ 5 Conclusions
Appendix A Discrepancy between $\sqrt{n}\sigma_{t}$ and the Distance $\text{dist}_{\mathcal{K}}(x_{t})$
Previous works Ho et al. (2020); Karras et al. (2022) have primarily focused on improving the estimation of the noise vector $\epsilon_{\theta}(x_{t},\sigma_{t})\approx\epsilon_{t}$ eq. 4, aligning with the first assumption of a zero-error denoiser. However, these approaches typically rely on a predefined noise level scheduler $\sigma_{t}$ for distance estimation, often without explicit validation. This reliance can lead to inaccuracies, even at the initial step, where $\text{dist}_{\mathcal{K}}(x_{T})\neq\sqrt{n}\sigma_{T}$.
@@ Appendix A Discrepancy between $\sqrt{n}\sigma_{t}$ and the Distance $\text{dist}_{\mathcal{K}}(x_{t})$
where the final equality uses the fact that $\mathbf{E}[|z_{T}|^{2}]=n$ for for $z_{T}\sim\mathcal{N}(0,I^{n\times n})$. Equation 22 implies that, with high probability, $\text{dist}_{\mathcal{K}}(x_{T})>\sqrt{n}\sigma_{T}$. Consequently, at any step $t=T,\dots,0$, these deviations may result in $\text{dist}_{\mathcal{K}}(x_{t})\neq\sqrt{n}\sigma_{t}$, potentially causing the final sample to deviate from the manifold $\mathcal{K}$. It is worth noting that the EDM sampling method initializes with a random step as $x_{T}=\sigma_{T}z_{T}$. However, it leads to the same conclusion: for $\langle z_{T},x_{0}^{\ast}\rangle\leq 0$, we have $\text{dist}_{\mathcal{K}}(x_{T})\neq\sqrt{n}\sigma_{T}$.
@@ Appendix B Sampling with Noise Level Correction
Algorithm 3 presents the algorithm for incorporating noise level correction within EDM. Unlike the approach used in DDIM with noise level correction Algorithm 1, the EDM version does not normalize the noise vector $\epsilon_{\theta}(x_{t},\hat{\sigma}_{t})$.
@@
0:   Noise scheduler $\sigma_{t}$
@@
Algorithm 4 presents the algorithm for incorporating noise level correction within DDNM for linear image restoration tasks $y=\textbf{A}x$.
@@
0:   Noise scheduler $\sigma_{t}$, randomness scale $\eta$,

### v3

@@ 3.3 Enhancing Sample Generation with Noise Level Correction
The trained noise level correction network $r_{\theta}$ can be integrated into various existing sampling algorithms to improve sample quality. For algorithms that include an initial sample estimate $\hat{x}_{0|t}$, eq. 5, we reformulate the one-step estimation as follows:
@@
We normalize the noise vector $\epsilon_{\theta}(\cdot)$ during the sampling as in eq. 17, process to decouple noise level (magnitude) correction $\hat{\sigma}_{t}$ from direction estimation $\hat{\epsilon}_{t}$. Empirical experiments show that normalizing $\epsilon_{\theta}(\cdot)$ and using $\hat{\sigma}_{t}$ to account for magnitude yields better results compared to not normalizing $\epsilon_{\theta}(\cdot)$. In the training loss function eq. 14, normalization of the noise vector $\epsilon$ is unnecessary because the randomly sampled noise $\epsilon$ naturally concentrates around its expected norm, approximately $\sqrt{n}$. However, the neural network-estimated noise vector $\epsilon_{\theta}(\cdot)$ does not necessarily maintain a constant norm.
@@
0:   Noise scheduler $\{\sigma_{t}\}_{t=1}^{T}$, randomness scale $\eta$ ($\eta=0$ for deterministic DDIM and $\eta=1$ for DDPM )
@@ 3.4 Constrained Sample Generation
Theorem 1 demonstrates that the projection operator $\text{proj}_{\mathcal{K}}(\cdot)$ can be approximately computed using the denoising diffusion models. Starting from an initial random point $x_{0}=\sigma_{\text{max}}\epsilon$, the projection onto the manifold $\mathcal{K}$ can be iteratively refined using eq. 16 and eq. 17, as follows:
@@
For the additional constraint projection, $\text{proj}_{\mathcal{C}}(x)$, the specific calculation depends on the nature of the constraint. The constraints for many image restoration tasks are linear, including inpainting, colorization, super-resolution, deblurring, and compressive sensing. For tasks with linear constraints, the projection can be computed directly or optimized using gradient descent. Consider an image restoration task formulated as $y=\textbf{A}x_{0}$, where $x_{0}$ represents the ground-truth image, $y$ is the degraded observation, A is the linear degradation operator. Given the degraded image $y$ and the current estimate $\hat{x}_{0|k}$, the projection onto the constraint can be computed as:
@@
Where $\textbf{A}^{\dagger}$ is the pseudo-inverse of A. In this work, we adopt the values of $\textbf{A}^{\dagger}$ for image restoration tasks as provided in (Wang et al., 2023). Here, ${x}_{0|k}$ is the $k-$th th estimate satisfying $x_{0}\in\mathcal{K}\cap\mathcal{C}$. This iterative estimation requires a predefined noise scheduler $\sigma_{1},\cdots,\sigma_{(k)},\cdots$ to generate ${x}_{0|k}$ and $x_{(k+1)}$. To allow flexible and potentially unlimited refinement steps, we define the noise schedule $\sigma_{(k)}$ with a maximum noise level $\sigma_{(0)}=\sigma_{max}$ and a minimum level $\sigma_{min}$, decaying by a predefined factor $\eta<1$. If the noise level reaches $\sigma_{min}$, the process can either stop with returning ${x}_{0|k}$ or restart from $\sigma_{restart}$. This strategy permits an arbitrary number of refinement steps, stopping either at a desired loss threshold or continuing indefinitely. Since $\sigma_{(k)}$ represents the distance of noisy samples from the manifold, this decaying schedule incrementally reduces $x_{(k)}$’s distance from the manifold.
@@ 4.1 Toy Experiments
We conducted a toy experiment to demonstrate the effectiveness of noise level correction in sample generation for diffusion models. The goal was to generate samples on $d$-sphere manifold. The toy training dataset was sampled from $d$-dimensional sphere manifold embedded within an $n$-dimensional data space, where $d<n$. Detailed experimental design information is available in Section C.1.
After training, we applied the proposed 10-step DDIM with Noise Level Correction (DDIM-NLC), as described in Algorithm 1, and compared it against the standard 10-step DDIM baseline. Sample quality was assessed by measuring the Euclidean distance from generated samples to the ground-truth manifold $\mathcal{K}$, with lower values indicating better quality. As discussed in Theorem 3.1, the noise level $\sqrt{n}\sigma_{t}$ approximates the distance to the manifold. To further evaluate this approximation, we examined the alignment between the estimated noise level $\hat{\sigma}_{t}$ and the true distance $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$.
Figure 4a shows the sample-to-manifold distances over time. The dashed curve indicates the predefined noise level $\sqrt{n}\sigma_{t}$, while the solid curves represent the measured distances from intermediate samples $x_{t}$ to the manifold. In the DDIM baseline, the actual distance deviates from $\sigma_{t}$, especially in later steps, highlighting the inaccuracy of predefined noise level assumptions (see also Figure 2). In contrast, the DDIM-NLC consistently produces samples that are closer to the manifold, demonstrating improved alignment through noise level correction. Figure 4b quantifies this improvement using the relative distance estimation bias, defined as:
@@
Where $\hat{\sigma}_{t}=\sigma_{t}$ for DDIM, and $\hat{\sigma}_{t}=\sigma_{t}[1+r_{\theta}(\hat{x}_{t},\sigma_{t})]$ for DDIM-NLC. Results show that DDIM-NLC significantly reduces estimation bias, especially in later steps when samples approach the manifold. These findings support the effectiveness of the proposed noise level correction strategy. Results for constrained sample generation can be found in Section C.3. The results show that DDIM-NLC achieves a significantly smaller distance estimation bias than DDIM, particularly in the later steps as samples approach the manifold. As illustrated in fig. 2, by correcting noise level estimation $\hat{\sigma_{t}}$, DDIM-NLC with more accurately estimates the distance. The results of constrained sample generation are shown in Section C.3.
(a) Comparison between $\text{dist}_{\mathcal{K}}(\hat{x}_{t})$ and $\sigma_{t}$
@@
Figure 4: Results from the toy experiment. (a) Sample-to-manifold distance over time for DDIM and DDIM-NLC. The dashed line shows the predefined noise level $\sqrt{n}\sigma_{t}$. (b) Relative bias in distance estimation.
@@ 4.2 Unconstrained Image Generation
We conducted experiments to evaluate the effectiveness of noise level correction in unconstrained image generation tasks. The noise level correction network was trained on top of a pre-trained denoiser network. Notably, the noise level correction network is approximately ten times smaller than the denoiser network. Additional details on the experimental setup are provided in Section D.1. We used the FID (Fréchet Inception Distance) score as the evaluation metric to assess the quality of generated samples, where lower scores indicate better quality, following standard practice in image generation tasks (Heusel et al., 2017). The experimental results for the DDIM/DDPM framework on the CIFAR-10 dataset are shown in Table 2. As observed, DDPM-NLC and DDIM-NLC, which incorporate noise level correction, outperform the original DDPM and DDIM models across all sampling steps. Specifically, our proposed noise level correction approach improves DDIM performance by 32%, 31%, 22%, 33%, and 28% for 10, 20, 50, 100, and 300 sampling steps, respectively.
@@ 4.3 Image Restoration
In this section, we evaluate the effectiveness of noise level correction on five common image restoration tasks, 4×4$\times$4 × super-resolution (SR) using bicubic downsampling, deblurring with a Gaussian blur kernel, colorization using an average grayscale operator, compressed sensing (CS) with a Walsh-Hadamard sampling matrix at a 0.25 compression ratio, and inpainting with text masks. These experiments are conducted on the ImageNet (Deng et al., 2009) and CelebA-HQ (Karras et al., 2018) datasets. We compare our method with recent state-of-the-art diffusion-based image restoration methods, including ILVR (Choi et al., 2021), RePaint (Lugmayr et al., 2022), DDRM (Kawar et al., 2022), and DDNM (Wang et al., 2023). For a fair comparison, all diffusion-based methods utilize the same pretrained denoising networks with the same 100-step denoising process (100 number of inference steps), following the experimental setup in (Wang et al., 2023). To evaluate sample quality, we use FID, PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index Measure). For colorization, where PSNR and SSIM are less effective metrics (Wang et al., 2023), we additionally use a Consistency metric, denoted as "Cons" and calculated as $\|Ax_{0}-y\|_{1}$. As a baseline, we also include the inverse solution for each image restoration task, given by $\hat{x}=\textbf{A}^{\dagger}y$, which achieves zero constraint violation but lacks the data manifold information.
@@ 5 Conclusions
Appendix A Insights for the Noise Level Correction
A.1 Discrepancy between $\sqrt{n}\sigma_{t}$ and the Distance $\text{dist}_{\mathcal{K}}(x_{t})$
Previous works (Ho et al., 2020; Karras et al., 2022) have primarily focused on improving the estimation of the noise vector $\epsilon_{\theta}(x_{t},\sigma_{t})\approx\epsilon_{t}$ eq. 4, aligning with the first assumption of a zero-error denoiser. However, these approaches typically rely on a predefined noise level scheduler $\sigma_{t}$ for distance estimation, often without explicit validation. This reliance can lead to inaccuracies, even at the initial step, where $\text{dist}_{\mathcal{K}}(x_{T})\neq\sqrt{n}\sigma_{T}$.
@@ A.1 Discrepancy between $\sqrt{n}\sigma_{t}$ and the Distance $\text{dist}_{\mathcal{K}}(x_{t})$
where the final equality uses the fact that $\mathbf{E}[|z_{T}|^{2}]=n$ for for $z_{T}\sim\mathcal{N}(0,I^{n\times n})$. Equation 22 implies that, with high probability, $\text{dist}_{\mathcal{K}}(x_{T})>\sqrt{n}\sigma_{T}$. Consequently, at any step $t=T,\dots,0$, these deviations may result in $\text{dist}_{\mathcal{K}}(x_{t})\neq\sqrt{n}\sigma_{t}$, potentially causing the final sample to deviate from the manifold $\mathcal{K}$. These errors arise from imperfections in the denoiser and become particularly significant in the final steps. The toy experiments on the discrepancy between noise level and true distance to the manifold can be found in section C.2. It is worth noting that the EDM sampling method initializes with a random step as $x_{T}=\sigma_{T}z_{T}$. However, it leads to the same conclusion: for $\langle z_{T},x_{0}^{\ast}\rangle\leq 0$, we have $\text{dist}_{\mathcal{K}}(x_{T})\neq\sqrt{n}\sigma_{T}$.
A.2 Training Objective
To address the issue of inaccurate distance estimation caused by relying on a predefined noise level $\sigma_{t}$, we propose a method called noise level correction (NLC) to better align the estimated noise level $\hat{\sigma}_{t}$ with the true distance $\text{dist}_{\mathcal{K}}(x_{t})/\sqrt{n}$. Specifically, we train a neural network to minimize the following objective (initial form):
$L:=\mathbb{E}\left[\left(\sqrt{n}\hat{\sigma}_{t}-\text{dist}_{\mathcal{K}}(x_{t})\right)^{2}\right]$ (23)
Inspired by residual learning, instead of directly predicting the distance, we propose to learn the residual between the true distance and the predefined noise level. We define:
$\hat{\sigma}_{t}:=\sigma_{t}\left[1+r_{\theta}(x_{t},\sigma_{t})\right]\approx\frac{\text{dist}_{\mathcal{K}}(x_{t})}{\sqrt{n}},$ (24)
where $r_{\theta}(x_{t},\sigma_{t})$ is the residual correction network to be trained. This residual formulation is advantageous because $r_{\theta}(\cdot)$ is typically more stable across noise levels, while $\sigma_{t}$ may vary widely during large diffusion time steps.
Since computing the true distance $\text{dist}_{\mathcal{K}}(x_{t})$ is impractical, we approximate it by the distance between a noisy sample and its clean counterpart obtained from the forward diffusion process. Given $x_{t}=x_{0}+\sigma_{t}\epsilon$, with $x_{0}\in\mathcal{K}$ and $\epsilon\sim\mathcal{N}(0,I)$, we approximate:
$\text{dist}_{\mathcal{K}}(x_{t})\approx\|x_{t}-x_{0}\|.$ (25)
This approximation is reasonable under the manifold hypothesis, assuming that the noise vector $\epsilon$ is orthogonal to the manifold, and that $\text{proj}_{\mathcal{K}}(x_{t})\approx x_{0}$. To further improve robustness, we introduce a perturbation scaling factor $\lambda$ and expand the input-output domain of $r_{\theta}(\cdot)$ via:
$x_{t}=x_{0}+\sigma_{t}\lambda\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),\quad\lambda\sim\mathcal{U}(1-\delta,1+\delta),$ (26)
where $\delta$ controls the perturbation range (e.g., $\delta=1$ recovers the original diffusion process). This leads to the approximation:
$\text{dist}_{\mathcal{K}}(x_{t})\approx\|x_{t}-x_{0}\|=\sigma_{t}\lambda\|\epsilon\|.$ (27)
Combining eq. 23, eq. 24, and eq. 27, we arrive at the final training objective, as shown in eq. 14:
$L_{r_{\theta}}:=\mathbf{E}_{x_{0},t,\epsilon,\lambda}\left[(\sqrt{n}\sigma_{t}[1+r_{\theta}(\hat{x}_{t},\sigma_{t})]-\sigma_{t}\lambda\|\epsilon\|)^{2}\right]$ (28)
$\hat{x}_{t}=x_{0}+\sigma_{t}\lambda\epsilon,~{}\epsilon\sim\mathcal{N}(0,I),~{}\lambda\sim\mathcal{U}(1-\delta,1+\delta)$ (29)
During training, we sample $x_{0}$ from the training set, $t$ from a uniform diffusion time range (e.g., $t\in[1,1000]$), and draw $\epsilon\sim\mathcal{N}(0,I)$ and $\lambda\sim\mathcal{U}(1-\delta,1+\delta)$. The loss is computed according to eq. 14 and used to optimize the noise level correction network $r_{\theta}(\cdot)$.
@@ Appendix B Sampling with Noise Level Correction
The proposed noise level correction (NLC) method is broadly compatible with a wide range of existing diffusion sampling algorithms and can be seamlessly integrated to enhance sample quality. In many diffusion samplers, each denoising step consists of two substeps: (1) estimating the original clean sample $\hat{x}_{0|t}$ given a noisy observation $x_{t}$, and (2) adding noise to propagate the sample to the next timestep. Our NLC method fits naturally into this structure by correcting the noise level during the clean sample estimation step, as described in eqs. 16 and 17. In sections 3.3 and 3.4, we demonstrated the integration of NLC with DDIM/DDPM and constrained sampling tasks. In this section, we extend our discussion to other samplers and show how the NLC method can be incorporated. All modifications to the original algorithms are highlighted in blue.
Algorithm 3 shows how NLC is integrated into the EDM sampler (Karras et al., 2022). Unlike DDIM-NLC in algorithm 1, EDM-NLC does not normalize the noise vector $\epsilon_{\theta}(x_{t},\hat{\sigma}_{t})$, since EDM is based on a second-order ODE solver. Nevertheless, the NLC method can be applied by modifying the noise level estimation step, as shown in blue.
@@
0:   Noise scheduler $\{\sigma_{t}\}_{t=1}^{T}$
@@
Algorithm 4 presents the incorporation of NLC into DDNM, a constrained image generation algorithm for linear inverse problems of the form $y=\mathbf{A}x$ (Wang et al., 2023). This demonstrates that the NLC method can also enhance constrained sampling tasks.
@@
0:   Noise scheduler $\{\sigma_{t}\}_{t=1}^{T}$, randomness scale $\eta$,
@@
Furthermore, the NLC method can be extended to other advanced samplers, such as DPM-Solver (ODE solver for Diffusion probabilistic models) (Lu et al., 2022). In algorithm 5, we present the integration of NLC into the second-order DPM-Solver. As with EDM-NLC, the primary modification involves adjusting the noise level using the NLC module. The DPM-Solver uses noise schedules $\{\alpha_{t}\}_{t=1}^{T}$ and $\{\sigma_{t}\}_{t=1}^{T}$, and defines the auxiliary parameter $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$. The inverse mapping from $\lambda$ to $t$ is denoted by $t_{\lambda}(\cdot)$, and we define the interval $h_{t}:=\lambda_{t}-\lambda_{t-1}$. These examples collectively demonstrate the flexibility of the proposed noise level correction strategy, making it a broadly applicable enhancement across various diffusion model sampling frameworks.
Algorithm 5 DPM-Solver with Noise Level Correction (DPM-NLC)
0:  Denoiser $\epsilon_{\theta}$ and noise level corrector $r_{\theta}$
0:   Noise scheduler $\{\alpha_{t}\}_{t=1}^{T}$ and $\{\sigma_{t}\}_{t=1}^{T}$
0:  samples $x_{0}\in\mathcal{K}$
1:  $x_{T}=\sigma_{T}\epsilon,~{}\epsilon\sim\mathcal{N}(0,I),$
2:  for $t=T,T-1,\cdots,1$ do
3:     $\hat{\sigma}_{t}\leftarrow\sigma_{t}\cdot[1+r_{\theta}(x_{t},\sigma_{t})]$
4:     $s=t_{\lambda}\left(\frac{\lambda_{t+1}+\lambda_{t}}{2}\right)$
5:     $\hat{\sigma}_{s}=\hat{\sigma}_{t}\frac{\sigma_{s}}{\sigma_{t}}$
6:     $\hat{\epsilon}_{t}=\epsilon_{\theta}(x_{t},\hat{\sigma}_{t})$
7:     $u=\frac{\alpha_{s}}{\alpha_{t}}x_{t}-\hat{\sigma}_{s}\left(e^{\frac{h_{t}}{2}}-1\right)\hat{\epsilon}_{t}$
8:     $\hat{\epsilon}_{s}=\epsilon_{\theta}(u,\hat{\sigma}_{s})$
9:     $x_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}-\sigma_{t-1}\left(e^{h_{t}}-1\right)\hat{\epsilon}_{s}$
10:  end for